OpenStudy (anonymous):

Megan’s backyard has a small pond that is 15 feet by 10 feet. Her mother wants her to add a wooden deck pathway around the pond, so that the total area of the pool and deck is 234 ft2. How wide does the pathway need to be to fulfill Megan’s mom’s request?

OpenStudy (anonymous):

 consider the length and breadth of the total portion be 15+x and 10+x resp... so the area is 234=(10+x)(15+x) by solving u will find x=3..hence the pathway shud be 3ft wide..